# The dawn is quiet II: Gaia XP constraints on the Milky Way’s proto-Galaxy from very metal-poor MDF tails

### v1

@@ Abstract
The earliest phase of the Milky Way’s evolution involved a transition from a dispersion-supported proto-galaxy to a rotationally supported disk. A key chemical signature of this transition is the moderate rise in $\alpha$/Fe] near $\mathrm{[Fe/H]}\approx-1.3$, which we previously interpreted as evidence for $\alpha$-enhanced gas accretion fueling early disk formation. However, this trend alone does not uniquely constrain the trade-off between initial gas mass, inflow rate, and star formation efficiency (SFE), leaving the physical condition of the proto-Milky Way uncertain. To break this degeneracy, we analyze the metal-poor tail ($3<\mathrm{[Fe/H]}<-2$) of the Milky Way’s metallicity distribution function (MDF) using Gaia DR3 BP/RP (XP) metallicities from ten catalogs. After applying recommended quality cuts, all catalogs exhibit a single-slope exponential tail with slopes $k\simeq 0.5$–2.0. Comparison with one-zone galactic chemical-evolution (GCE) models that replicated the $\alpha$/Fe]-rise from Paper I shows that shallow tails ($k\simeq 0.6$) require a massive initial cold gas reservoir ($\gtrsim 10^{9}\,\mathrm{M_{\odot}}$), while steeper tails ($k\gtrsim 1$) arise from small reservoirs that built up over time with weak inflow. MDFs with $k\simeq 1.0$ are best reproduced under our GCE framework, which favor a proto-Galaxy with a moderate gas reservoir ($10^{8}$$10^{9}\,\mathrm{M_{\odot}}$) sustained through weak continuous inflow ($\sim 2\ \mathrm{M_{\odot}\ yr^{-1}}$) and SFE comparable to today’s value (a few $\times 10^{-10}\,\mathrm{yr^{-1}}$) during the first Gyr. This scenario is reinforced by MDFs of 30 Milky Way analogs in the Auriga simulations, which exhibit similar slopes ($k\approx 1.2$). The metal-poor MDF tail thus provides a quantitative constraint on the Milky Way’s early gas accretion and star formation history.
@@ 1 Introduction
The earliest phase of the Milky Way’s formation was dominated by a rapid and turbulent phase in the form of a compact proto-galaxy, preceding the emergence of the rotationally supported disk. Elemental abundances from APOGEE and kinematics from Gaia indicate an in-situ, kinematically hot “Aurora” phase at low metallicity, followed by a fast “spin-up” of the Galaxy as its disk started to form between $1.3<[\mathrm{Fe/H}]<-0.9$ over $\sim$1–2 Gyr, with a simultaneous drop in abundance scatter (2022MNRAS.514..689B). A central concentration of 18,000 giant stars with [Fe/H] <1.5<-1.5 have been found within $\sim 5$ kpc of the Galactic center and proposed to be the proto-Galactic nucleus. These stars suggest that the metal-poor “heart” contains $10^{8}\ \mathrm{M_{\odot}}$ and formed >12.5>12.5 Gyr ago (2022ApJ...941...45R). Multiple components belonging to the proto-Galaxy have also been identified and characterized with a mix of in situ vs accreted origins (2024ApJ...964..104M; 2025MNRAS.537.3730H).
@@
The disk of the Milky Way commenced its assembly around 13 Gyr ago, with its star formation rate (SFR) peaking around 11 Gyr ago, coinciding with the Gaia-Sausage-Enceladus (GSE) (2022Natur.603..599X; 2021MNRAS.508.5903V; 2022MNRAS.514..689B; 2025NatAs...9..101X). This used to be considered unusually early, but a cold rotating disk has been confirmed by an extra-galactic observation at $z\sim 4.3$ when the universe was only 1.5 Gyr old (2020Natur.581..269N). The quiescent evolution of the disk took over around 8 Gyr ago and continued to the present as evidenced by a clear division in the stellar age-metallicity distribution (2013A&A...560A.109H; 2015A&A...578A..87S; 2019MNRAS.489.1742F; 2022MNRAS.510.4669S). These three phases of evolution are confirmed by a sample of ten million giants from Gaia (2024ApJ...972..112C). However, there is evidence that the kinematically defined thin and thick disk could have also co-evolved from 13 Gyr ago (2024A&A...688A.167N; 2025arXiv250400135B).
@@
As the Galactic disk began to form, a moderate rise in $\alpha$/Fe] appeared near $\mathrm{[Fe/H]}\approx-1.3$, even among stars on low-eccentricity orbits (2022arXiv220402989C; 2022MNRAS.514..689B; 2022ApJ...941...45R). This feature marks the transition from the turbulent, kinematically hot Aurora phase to the onset of ordered, rotationally supported star formation and has been interpreted as the chemical signature of the disk’s birth. In Paper I (2024PASA...41...63C), we showed that the observed $\alpha$/Fe] upturn is best explained by the sudden onset of $\alpha$-enhanced gas inflow-pre-enriched by core-collapse supernovae in the halo-which supplied both angular momentum and fuel for renewed star formation after a preceding inflow-poor phase. The resulting enhancement thus reflects the coupled chemical and dynamical effects of the accreted material. However, the $\alpha$/Fe]-rise alone cannot uniquely determine the proto-Galaxy’s initial gas mass, star-formation efficiency, or inflow timescale, as multiple parameter combinations yield similar patterns. To break this degeneracy, we now turn to the slope of the metal-poor tail of the metallicity distribution function (MDF), which traces the balance between gas supply and metal production during the Galaxy’s earliest evolution.
The all-sky BP/RP spectrophotometry (XP spectra) from Gaia DR3 offers an unprecedented opportunity to extend the analysis beyond the $\alpha$/Fe] feature and directly probe the metal-poor tail of the Galactic MDF. Because the XP selection is essentially photometric and astrometric-without pre-selection on kinematics, latitude, or metallicity-it provides a nearly unbiased census of stars across the Milky Way. Covering 330-1050 nm for about 220 million sources observed with a single, homogeneous instrument and calibration pipeline, XP spectra capture strong metal-sensitive features such as Ca ii H & K. When combined with parallaxes and broadband photometry, these data break key degeneracies among stellar parameters, enabling uniform metallicity estimates that reach into the Galactic plane and low-$|\mathrm{z}|$ regions long under-sampled by spectroscopic surveys. Building on Paper I, we now exploit the XP-based metallicity catalogs to study the logarithmic MDF over $3<\mathrm{[Fe/H]}<-2.0$. Using a grid of one-zone GCE models that replicated the $\alpha$/Fe]-rise in Paper I, we interpret the observed slopes to constrain the proto-Galaxy’s parameters.
@@
Very metal-poor (VMP; $\mathrm{[Fe/H]}<-2.0$) and extremely metal-poor (EMP; $\mathrm{[Fe/H]}<-3.0$) stars provide the most direct fossil record of the proto-Galaxy, retaining the chemical and kinematic signatures of the first star-formation episodes. Even small, well-characterized samples offer strong constraints on early enrichment and gas accretion-for instance, asteroseismic ages of $\sim$ 12 Gyr for stars with $\mathrm{[Fe/H]}\approx-2.7$ set lower bounds on the onset of in-situ disk formation (2024ApJ...975...19H). Interpreting such ancient populations requires GCE modeling, which links observables such as the MDF, $\alpha$/Fe]-[Fe/H] trends, and age-abundance relations to the underlying inflow, outflow, and star formation histories (SFH). Historically, VMP and EMP stars have constrained the critical metallicity for low-mass star formation and the yields of the first supernovae (1976ApJ...209..418H; 2003A&A...404..211P; 2007MNRAS.381..647S; 2013A&A...554A.135B), but robust statistical inference has been limited by the rarity of such stars and strong selection biases. Gaia XP now provides the first opportunity to measure their metallicity distribution function homogeneously across the Galaxy, enabling us to quantify the metal-poor tail and use it to infer the proto-Milky Way’s earliest gas-regulation conditions. We restrict our analysis to $\mathrm{[Fe/H]}>-3.0$, as metallicity estimates below are less reliable and such extremely low metallicities are typically skipped in one or two steps in GCE models, requiring dedicated modeling beyond the scope of this work.
Cosmological simulations provide a broader framework for the evolution of disk galaxies. Milky Way-mass progenitors at $z\gtrsim 2$ undergo bursty star formation fueled by cold, dense gas inflows and regulated by energetic stellar feedback (2006MNRAS.368....2D; 2015MNRAS.454.2691M; 2021ApJ...911...88S). Their proto-galactic components are typically compact, with half-mass radii of 5–10 kpc, and form early (8–13 Gyr ago; $z\sim$ 1–6) with only mild net rotation and $\alpha$-enhanced, metal-poor chemistry resembling that of the oldest Galactic disk stars. The proto-galaxies are usually dominated by one or two major progenitors plus a few smaller companions (2024MNRAS.527.9810H). Although centrally concentrated, their oldest stars are dispersed across the halo by repeated starburst and feedback episodes common at $z\gtrsim 2$ (2018MNRAS.480..652E). As inflow stabilizes, a coherent stellar disk can arise within the first $\sim$ 1–2 Gyr of cosmic time, forming rapidly (within $\sim$ 0.2 Gyr) amid residual bursts and surviving strong feedback, ultimately achieving the rotation and $\alpha$-enhanced chemistry consistent with the Milky Way’s most metal-poor stellar populations (2023MNRAS.519.2598G; 2025ApJ...990....7S).
@@
In this paper, we use Gaia DR3 XP metallicities to quantify the metal-poor tail of the Milky Way’s metallicity distribution function (MDF) over $3<\mathrm{[Fe/H]}<-2.0$ and interpret the measured slopes within a one-zone GCE framework. Building on Paper I, which linked the observed rise in $\alpha$/Fe] at $\mathrm{[Fe/H]}\approx-1.3$ to the onset of $\alpha$-enhanced inflow and disk spin-up, we now constrain the proto-Galaxy’s parameters using a complementary diagnostic: the logarithmic slope of the metal-poor MDF tail. This approach leverages Gaia XP’s homogeneous, Galaxy-wide sampling of VMP stars to test whether the Milky Way’s earliest phase resembled the burst-regulated, gas-rich progenitors predicted by cosmological simulations. Together, these analyses aim to provide a quantitative, physically motivated description of the conditions under which the Galactic disk first emerged.
2 Gaia XP metallicity catalogs
@@
Before Gaia DR3, assembling large, homogeneous samples of very metal-poor stars relied on surveys that either used low-resolution objective-prism spectroscopy (HK, HES) or medium-resolution spectroscopic campaigns (SEGUE, RAVE). Photometric programs such as SkyMapper and Pristine employed metallicity-sensitive filters for efficient pre-selection, while LAMOST provided extensive low-resolution spectroscopy from which metal-poor candidates could be drawn. Compiled databases like SAGA and JINAbase aggregated high-resolution measurements but inherited heterogeneous selection functions. These efforts established the main strategies used to identify metal-poor stars—low-resolution spectral indicators of line blanketing, photometric metallicity proxies, and broad spectroscopic classification—and they set the historical context for interpreting Gaia XP data.
@@
The Gaia XP spectra, released in DR3, provide low-resolution ($R\sim 30$100100) spectrophotometry from the BP and RP instruments covering 330–1050 nm for sources down to $G\approx 17.65$. Although the resolving power is significantly lower than that of most spectroscopic surveys, the all-sky coverage, homogeneous calibration, and continuous spectral sampling encode metallicity information for tens of millions of stars, including very metal-poor populations. XP shares methodological features with earlier objective-prism and photometric surveys—both rely on broad-band spectral energy distributions and weak metal-line signatures—but with the advantage of contiguous spectrophotometric data and a uniform instrument response.
@@
The Gaia DR3 pipeline serves as the baseline of analyzing the XP spectra: GSP-Phot fits XP spectra plus parallax and $G$ to synthetic libraries in a Bayesian forward model, producing $T_{\mathrm{eff}}$, $\log\,g$, and a calibrated metallicity for $\sim 471\,\mathrm{M}$ sources (useful mostly qualitatively for $\mathrm{M}/\mathrm{H}]$ per their own validation) (2023A&A...674A..27A). The signal-to-noise ratio (S/N) of the XP coefficients is a key factor in the reliability of these estimates. Building on that, 2023ApJS..267....8A trained gradient-boosted trees on APOGEE DR17 labels augmented with a curated very metal-poor set, and fed the model XP, Gaia, broadband, narrowband (including synthetic Pristine CaHK) and CatWISE photometry as features; they predict $T_{\mathrm{eff}}$, $\log\,g$, $\mathrm{M}/\mathrm{H}]$ for $\sim 175\,\mathrm{M}$ stars and also provide a carefully vetted bright RGB subset ($G\lesssim 16$) where the metallicity is most reliable.
@@
SHBoost (2024A&A...691A..98K) targets essentially “all stars with XP and parallaxes”: it learns $T_{\mathrm{eff}}$, $\log\,g$, $\mathrm{M}/\mathrm{H}]$, $A_{V}$, and mass with XGBoost trained on StarHorse labels compiled from many spectroscopic surveys (plus curated hot/VMP/WD sub-samples) (2018MNRAS.476.2556Q) using XP coefficients together with Gaia+2MASS+WISE; the team also applies a posteriori calibration for metallicity systematics. 2023MNRAS.524.1855Z go the other way—forward-modelling XP and broadband photometry with a data-driven model trained on LAMOST; they provide $T_{\mathrm{eff}}$, $\log\,g$, $\mathrm{Fe}/\mathrm{H}]$ (and revised distances/extinctions) for $\sim 220\,\mathrm{M}$ stars, i.e., an all-sky, all-types sample anchored to LAMOST labels.
@@
Several recent XP-based analyses apply more targeted strategies, either by improving the XP flux calibration before modeling or by restricting the training domain to regimes where the abundance information is strongest. 2025A&A...695A..75Y first infer per-star flux corrections with an APOGEE-trained neural network to suppress small-scale XP wiggles and then fit the corrected spectra with FERRE against synthetic grids, producing a $68\,\mathrm{M}$-star catalog of $T_{\mathrm{eff}}$, $\log\,g$, and $\mathrm{M}/\mathrm{H}]$ with enhanced control of systematics across the sky. AspGap (2024ApJS..272....2L) focuses on red-giant branch stars and employs a neural network trained on APOGEE that includes an auxiliary component designed to reconstruct the corresponding high-resolution APOGEE spectrum. This additional reconstruction objective strengthens the link between XP features and abundance-sensitive spectral structure, yielding $T_{\mathrm{eff}}$, $\log\,g$, $\mathrm{M}/\mathrm{H}]$, and precise $\alpha/\mathrm{M}]$ for $\sim 37\,\mathrm{M}$ RGB stars.
@@
2025ApJ...980...90H adopt a complementary strategy optimized for low-extinction sightlines. They apply explicit cuts in Galactic latitude and E(BV)E(B-V), impose a $G<18.85$ limit, and train boosted trees on APOGEE DR17 augmented with a curated very metal-poor sample, publishing $\mathrm{M}/\mathrm{H}]$ and $\alpha/\mathrm{M}]$ for $\sim 48\,\mathrm{M}$ stars in these dust-minimal windows. 2025ApJS..279....7Y emphasize giants as well, dereddening and correcting XP spectra before using an uncertainty-aware, cost-sensitive neural network (UA-CSNet) trained on high-quality spectroscopic labels (PASTEL) to predict $\mathrm{Fe}/\mathrm{H}]$ for $\sim 20\,\mathrm{M}$ giants with particular attention to the metal-poor tail.
@@
Two efforts push specific elemental abundances beyond metallicity. 2025MNRAS.537.1984A train a neural network on spectroscopic reference labels to infer both $\mathrm{Fe}/\mathrm{H}]$ and $\mathrm{C}/\mathrm{Fe}]$ from XP (plus auxiliary Gaia information) for $10\,\mathrm{M}$ stars, then identify $\sim 2{,}000$ bright carbon-enhanced metal-poor (CEMP) candidates; their tables link back to vetted giant selections to ensure reliability at the metal-poor end. 2025MNRAS.536.2507K focus on abundance patterns associated with “second-generation” globular cluster (GC) populations. They employ a multilayer perceptron (MLP) neural network trained on APOGEE cross-matches to predict $T_{\mathrm{eff}}$, $\log\,g$, $\mathrm{Fe}/\mathrm{H}]$, $\mathrm{N}/\mathrm{O}]$, and $\mathrm{Al}/\mathrm{Fe}]$ directly from normalized XP coefficients. Applied to the Andrae-vetted RGB sample, these predictions allow them to identify high-$\mathrm{N}/\mathrm{O}]$ field-star candidates that may trace a GC-origin channel. Finally, one catalog uses XP to synthesize a metallicity-sensitive narrow band: the Pristine DR1-XP release computes synthetic CaHK magnitudes for $\sim 219\,\mathrm{M}$ stars and, combined with broadband Gaia (and deeper Pristine where available), publishes photometric metallicities for $30\,\mathrm{M}$ high-S/N FGK stars, with strong performance below $\mathrm{Fe}/\mathrm{H}]\approx-1$ and extending to $\lesssim-3.5$ (2024A&A...692A.115M).

### v2

@@ Abstract
The earliest phase of the Milky Way’s evolution involved a transition from a dispersion-supported proto-galaxy to a rotationally supported disk. A key chemical signature of this transition is the moderate rise in $\alpha$/Fe] near $\mathrm{[Fe/H]}\approx-1.3$, which we previously interpreted as evidence for $\alpha$-enhanced gas accretion fueling early disk formation. However, this trend alone does not uniquely constrain the trade-off between initial gas mass, inflow rate, and star formation efficiency (SFE), leaving the physical condition of the proto-Milky Way uncertain. To break this degeneracy, we analyze the metal-poor tail ($3<\mathrm{[Fe/H]}<-2$) of the Milky Way’s metallicity distribution function (MDF) using Gaia DR3 BP/RP (XP) metallicities from ten catalogs. After applying recommended quality cuts, all catalogs exhibit a single-slope exponential tail with slopes $k\simeq 0.6$–2.0. Comparison with one-zone galactic chemical-evolution (GCE) models that replicated the $\alpha$/Fe]-rise from Paper I shows that shallow tails ($k\simeq 0.6$) require a massive initial cold gas reservoir ($\gtrsim 10^{9}\,\mathrm{M_{\odot}}$), while steeper tails ($k\gtrsim 1$) arise from small reservoirs that built up over time with weak inflow. MDFs with $k\simeq 1.0$ are best reproduced under our GCE framework, which favor a proto-Galaxy with a moderate gas reservoir ($10^{8}$$10^{9}\,\mathrm{M_{\odot}}$) sustained through weak continuous inflow ($\sim 2\ \mathrm{M_{\odot}\ yr^{-1}}$) and SFE comparable to today’s value (a few $\times 10^{-10}\,\mathrm{yr^{-1}}$) during the first Gyr. This scenario is reinforced by MDFs of 30 Milky Way analogs in the Auriga simulations, which exhibit similar slopes ($k\approx 1.25$). The metal-poor MDF tail thus provides a quantitative constraint on the Milky Way’s early gas accretion and star formation history.
@@ I Introduction
The earliest phase of the Milky Way’s formation was dominated by a rapid and turbulent phase in the form of a compact proto-galaxy, preceding the emergence of the rotationally supported disk. Elemental abundances from APOGEE (Majewski et al., 2017) and kinematics from Gaia (Gaia Collaboration et al., 2016) indicate an in-situ, kinematically hot “Aurora” phase at low metallicity, followed by a fast “spin-up” of the Galaxy as its disk started to form between $1.3<[\mathrm{Fe/H}]<-0.9$ over $\sim$1–2 Gyr, with a simultaneous drop in abundance scatter (Belokurov & Kravtsov, 2022). A central concentration of 18,000 giant stars with [Fe/H] <1.5<-1.5 have been found within $\sim 5$ kpc of the Galactic center and proposed to be the proto-Galactic nucleus. These stars suggest that the metal-poor “heart” contains $10^{8}\ \mathrm{M_{\odot}}$ and formed >12.5>12.5 Gyr ago (Rix et al., 2022). Multiple components belonging to the proto-Galaxy have also been identified and characterized with a mix of in situ vs accreted origins (Malhan & Rix, 2024; Horta & Schiavon, 2025).
@@
The disk of the Milky Way commenced its assembly around 13 Gyr ago, with its star formation rate (SFR) peaking around 11 Gyr ago, coinciding with the Gaia-Sausage-Enceladus (GSE) (Xiang & Rix, 2022; Vincenzo et al., 2021; Belokurov & Kravtsov, 2022; Xiang et al., 2025). This used to be considered unusually early, but a cold rotating disk has been confirmed by an extra-galactic observation at $z\sim 4.3$ when the universe was only 1.5 Gyr old (Neeleman et al., 2020). The quiescent evolution of the disk took over around 8 Gyr ago and continued to the present as evidenced by a clear division in the stellar age-metallicity distribution (Haywood et al., 2013; Snaith et al., 2015; Feuillet et al., 2019; Sahlholdt et al., 2022). These three phases of evolution are confirmed by a sample of ten million giants from Gaia (Chandra et al., 2024). However, there is evidence that the kinematically defined thin and thick disk could have also co-evolved from 13 Gyr ago (Nepal et al., 2024; Borbolato et al., 2025).
@@
As the Galactic disk began to form, a moderate rise in $\alpha$/Fe] appeared near $\mathrm{[Fe/H]}\approx-1.3$, even among stars on low-eccentricity orbits (Conroy et al., 2022; Belokurov & Kravtsov, 2022; Rix et al., 2022). This feature marks the transition from the turbulent, kinematically hot Aurora phase to the onset of ordered, rotationally supported star formation and has been interpreted as the chemical signature of the disk’s birth. In Paper I (Chen et al., 2024), we showed that the observed $\alpha$/Fe] upturn is best explained by the sudden onset of $\alpha$-enhanced gas inflow-pre-enriched by core-collapse supernovae in the halo-which supplied both angular momentum and fuel for renewed star formation after a preceding inflow-poor phase. The resulting enhancement thus reflects the coupled chemical and dynamical effects of the accreted material. However, the $\alpha$/Fe]-rise alone cannot uniquely determine the proto-Galaxy’s initial gas mass, star-formation efficiency, or inflow timescale, as multiple parameter combinations yield similar patterns. To break this degeneracy, we now turn to the slope of the metal-poor tail of the metallicity distribution function (MDF), which traces the balance between gas supply and metal production during the Galaxy’s earliest evolution.
@@
Very metal-poor (VMP; $\mathrm{[Fe/H]}<-2.0$) and extremely metal-poor (EMP; $\mathrm{[Fe/H]}<-3.0$) stars provide the most abundant fossil record of the proto-Galaxy, retaining the chemical and kinematic signatures of the first star-formation episodes. Even small, well-characterized samples offer strong constraints on early enrichment and gas accretion-for instance, asteroseismic ages of $\sim$ 12 Gyr for stars with $\mathrm{[Fe/H]}\approx-2.7$ set lower bounds on the onset of in-situ disk formation (Huber et al., 2024). Interpreting such ancient populations requires GCE modeling, which links observables such as the MDF, $\alpha$/Fe]-[Fe/H] trends, and age-abundance relations to the underlying inflow, outflow, and star formation histories (SFH). Historically, VMP and EMP stars have constrained the critical metallicity for low-mass star formation and the yields of the first supernovae (Hartwick, 1976; Prantzos, 2003; Salvadori et al., 2007; Brusadin et al., 2013), but robust statistical inference has been limited by the rarity of such stars and strong selection biases. Gaia DR3 BP/RP spectrophotometry (XP spectra) (Gaia Collaboration et al., 2023; De Angeli et al., 2023) now makes it possible to measure the metal-poor tail of the Galactic MDF with homogeneous, all-sky stellar-parameter catalogs far larger than any previous spectroscopic sample.
[…]
Although early star formation is centrally concentrated, repeated burst–feedback cycles can dynamically heat stellar orbits and disperse the oldest populations into spheroidal, halo-like configurations (El-Badry et al., 2016, 2018). The resulting proto-galaxies are typically built by one or two dominant progenitors plus a small number of lower-mass companions (Horta et al., 2024). In Auriga, the early stellar component remains compact and metal-poor, with mild net rotation and $\alpha$-enhanced chemistry broadly resembling the oldest Galactic disk populations, while continued cosmological inflow sustains star formation and generates extended metal-poor populations (Grand et al., 2017). The transition to an ordered, rotationally supported stellar disk occurs as inflow and feedback-driven turbulence become less disruptive due to the change in the gravitational potential. Once gas accretion stabilizes, a coherent disk can emerge within the first $\sim$1–2 Gyr of cosmic time and, in some cases, assemble rapidly (within $\sim$0.2 Gyr) despite residual burstiness, attaining the rotation and $\alpha$-enhanced chemical patterns characteristic of the Milky Way’s most metal-poor disk stars (Yu et al., 2021; Gurvich et al., 2023; Hopkins et al., 2023; Semenov et al., 2025).
@@
In this paper, we use Gaia DR3 XP metallicities to quantify the metal-poor tail of the Milky Way’s MDF over $3<\mathrm{[Fe/H]}<-2.0$ and interpret the measured slopes within a one-zone GCE framework. Building on Paper I, which linked the observed rise in $\alpha$/Fe] at $\mathrm{[Fe/H]}\approx-1.3$ to the onset of $\alpha$-enhanced inflow and disk spin-up, we now constrain the proto-Galaxy’s parameters using a complementary diagnostic: the logarithmic slope of the metal-poor MDF tail. This approach leverages Gaia XP’s homogeneous, Galaxy-wide sampling of VMP stars to test whether the Milky Way’s earliest phase resembled the burst-regulated, gas-rich progenitors predicted by cosmological simulations. Together, these analyses aim to provide a quantitative, physically motivated description of the conditions under which the Galactic disk first emerged.
This paper is organized as follows. Section II summarizes the Gaia DR3 XP metallicity catalogs used in this work and the adopted quality cuts, and presents basic CMD and spatial/kinematic diagnostics to characterize their selection footprints and potential contamination. Section III describes the GCE model grid (introduced in Paper I) and the procedure for predicting and measuring metal-poor MDF tail slopes. In Section IV we measure the logarithmic MDF slopes over $3<\mathrm{[Fe/H]}<-2.0$ across catalogs, test their robustness, and interpret the inferred $k$ values in terms of the proto-Galaxy’s parameters. We then compare the observed slopes to Milky Way analogs in the Auriga simulations. Section V discusses systematic uncertainties and the connection to prior halo MDF and GCE studies. Section VI summarizes our conclusions.
II Gaia XP metallicity catalogs
II.1 Before Gaia
Before Gaia DR3, assembling large and homogeneous samples of very metal-poor stars relied on a sequence of survey strategies that progressively expanded both sky coverage and statistical accuracy. Early objective-prism work demonstrated that the Ca II H&K resonance lines provide a robust metallicity diagnostic at low spectral resolution; in particular, Bessell & Norris (1984) identified a star with $\mathrm{Fe/H}]=-4.5$, establishing the feasibility of systematically probing the metal-poor tail. Subsequent discoveries pushed the metallicity frontier to the ultra-metal-poor regime ($\mathrm{[Fe/H]}<-4$) and revealed surprising findings on their elemental abundances (Christlieb et al., 2002; Frebel et al., 2005; Caffau et al., 2011).
@@ II.1 Before Gaia
This approach was scaled to large-area surveys such as the HK survey (Beers et al., 1985, 1992) and the Hamburg/ESO (HES) survey (Wisotzki et al., 2000; Christlieb et al., 2008), transforming isolated discoveries into statistically meaningful halo samples. Medium-resolution spectroscopic campaigns, including SEGUE (Yanny et al., 2009) and RAVE (Smith et al., 2007; Matijevič et al., 2017), further expanded the census of metal-poor stars across wide sky areas. Complementary photometric programs such as SkyMapper (Keller et al., 2007) and Pristine (Starkenburg et al., 2017) employed metallicity-sensitive filters for efficient candidate pre-selection, while LAMOST (Zhao et al., 2012) provided extensive low-resolution spectroscopy from which metal-poor targets could be drawn. Compiled databases such as SAGA (Suda et al., 2008) and JINAbase (Abohalima & Frebel, 2018) aggregated high-resolution abundance measurements, though with heterogeneous selection functions. Together, these efforts established the principal methodologies for identifying metal-poor stars—low-resolution spectral indicators of line blanketing, photometric metallicity proxies, and broad spectroscopic classification—and they form the historical foundation against which the homogeneous Gaia XP metallicity measurements can be interpreted.
II.2 Training Methods for Gaia Catalogs
@@
The Gaia XP spectra, released in DR3, provide low-resolution ($R\sim 20$$100$) spectrophotometry from the BP and RP instruments covering 330–1050 nm for sources down to $G\approx 17.65$ (De Angeli et al., 2023). Although the resolving power is significantly lower than that of most spectroscopic surveys, the all-sky coverage, homogeneous calibration, and continuous spectral sampling encode metallicity information for tens of millions of stars, including very metal-poor populations. XP shares methodological features with earlier objective-prism and photometric surveys—both rely on broad-band spectral energy distributions and weak metal-line signatures—but with the advantage of contiguous spectrophotometric data and a uniform instrument response.
@@
We provide a brief summary of the various techniques used to derive metallicities from Gaia XP spectra here. The Gaia DR3 pipeline serves as the baseline of analyzing the XP spectra (see GaiaXPy [^1] https://www.cosmos.esa.int/web/gaia/gaiaxpy for more details). GSP-Phot fits XP spectra plus parallax and $G$ to synthetic libraries in a Bayesian forward model, producing $T_{\mathrm{eff}}$, $\log\,g$, and a calibrated metallicity for $\sim 471\,\mathrm{M}$ sources (useful mostly qualitatively for $\mathrm{M}/\mathrm{H}]$ per their own validation) (Andrae et al., 2023b). The signal-to-noise (S/N) ratio of the XP coefficients is a key factor in the reliability of these estimates. Building on that, Andrae et al. (2023a) trained gradient-boosted trees on APOGEE DR17 labels augmented with a curated very metal-poor set, and fed the model XP, Gaia, broadband, narrowband (including synthetic Pristine CaHK) and CatWISE photometry (Marocco et al., 2021) as features; they predict $T_{\mathrm{eff}}$, $\log\,g$, $\mathrm{M}/\mathrm{H}]$ for $\sim 175\,\mathrm{M}$ stars and also provide a carefully vetted bright RGB subset ($G\lesssim 16$) where the metallicity is most reliable.
@@
SHBoost (Khalatyan et al., 2024) targets essentially “all stars with XP and parallaxes”: it learns $T_{\mathrm{eff}}$, $\log\,g$, $\mathrm{M}/\mathrm{H}]$, $A_{V}$, and mass with XGBoost trained on StarHorse labels compiled from many spectroscopic surveys (plus curated hot/VMP/WD sub-samples) (Queiroz et al., 2018) using XP coefficients together with Gaia+2MASS(Skrutskie et al., 2006)+WISE(Wright et al., 2010); the team also applies a posteriori calibration for metallicity systematics. Zhang et al. (2023) go the other way—forward-modelling XP and broadband photometry with a data-driven model trained on LAMOST; they provide $T_{\mathrm{eff}}$, $\log\,g$, $\mathrm{Fe}/\mathrm{H}]$ (and revised distances/extinctions) for $\sim 220\,\mathrm{M}$ stars, i.e., an all-sky, all-types sample anchored to LAMOST labels.
@@
Several recent XP-based analyses apply more targeted strategies, either by improving the XP flux calibration before modeling or by restricting the training domain to regimes where the abundance information is strongest. Ye et al. (2025) first infer per-star flux corrections with an APOGEE-trained neural network to suppress small-scale XP wiggles and then fit the corrected spectra with FERRE against synthetic grids, producing a $68\,\mathrm{M}$-star catalog of $T_{\mathrm{eff}}$, $\log\,g$, and $\mathrm{M}/\mathrm{H}]$ with enhanced control of systematics across the sky. AspGap (Li et al., 2024) focuses on red-giant branch stars and employs a neural network trained on APOGEE that includes an auxiliary component designed to reconstruct the corresponding high-resolution APOGEE spectrum. This additional reconstruction objective strengthens the link between XP features and abundance-sensitive spectral structure, yielding $T_{\mathrm{eff}}$, $\log\,g$, $\mathrm{M}/\mathrm{H}]$, and precise $\alpha/\mathrm{M}]$ for $\sim 37\,\mathrm{M}$ RGB stars.
@@
Hattori (2025) adopt a complementary strategy optimized for low-extinction sightlines. They apply explicit cuts in Galactic latitude and E(BV)E(B-V), impose a $G<18.85$ limit, and train boosted trees on APOGEE DR17 augmented with a curated very metal-poor sample, publishing $\mathrm{M}/\mathrm{H}]$ and $\alpha/\mathrm{M}]$ for $\sim 48\,\mathrm{M}$ stars in these dust-minimal windows. Yang et al. (2025) emphasize giants as well, dereddening and correcting XP spectra before using an uncertainty-aware, cost-sensitive neural network (UA-CSNet) trained on high-quality spectroscopic labels (PASTEL) to predict $\mathrm{Fe}/\mathrm{H}]$ for $\sim 20\,\mathrm{M}$ giants with particular attention to the metal-poor tail.
@@
Two efforts push specific elemental abundances beyond metallicity. Ardern-Arentsen et al. (2025) train a neural network on spectroscopic reference labels to infer both $\mathrm{Fe}/\mathrm{H}]$ and $\mathrm{C}/\mathrm{Fe}]$ from XP (plus auxiliary Gaia information) for $10\,\mathrm{M}$ stars, then identify $\sim 2{,}000$ bright carbon-enhanced metal-poor (CEMP) candidates; their tables link back to vetted giant selections to ensure reliability at the metal-poor end. Kane et al. (2025) focus on abundance patterns associated with “second-generation” globular cluster (GC) populations. They employ a multilayer perceptron (MLP) neural network trained on APOGEE cross-matches to predict $T_{\mathrm{eff}}$, $\log\,g$, $\mathrm{Fe}/\mathrm{H}]$, $\mathrm{N}/\mathrm{O}]$, and $\mathrm{Al}/\mathrm{Fe}]$ directly from normalized XP coefficients. Applied to the Andrae-vetted RGB sample, these predictions allow them to identify high-$\mathrm{N}/\mathrm{O}]$ field-star candidates that may trace a GC-origin channel. Finally, one catalog uses XP to synthesize a metallicity-sensitive narrow band: the Pristine DR1-XP release computes synthetic CaHK magnitudes for $\sim 219\,\mathrm{M}$ stars and, combined with broadband Gaia (and deeper Pristine where available), publishes photometric metallicities for $30\,\mathrm{M}$ high-S/N FGK stars, with strong performance below $\mathrm{Fe}/\mathrm{H}]\approx-1$ and extending to $\lesssim-3.5$ (Martin et al., 2024).